RC Meeting and Run Plan, Thursday, December 12th 2019

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Current Status

  • RC Daily Update [1]
  • Moller Commisioning currently underway
  • Ramp Lock done during DAY
  • New Moller Optics tested this morning. Spot size improved
  • Both spectrometers at 30 degrees/2GeV; SHMS configured with defocussed optics (Q2 set +20%)[2]

Current Run Plan

  • Daily Run Plan and Run Status
  • Priorities
    • First: Spin Dance until done (tomorrow)
    • Next: commission and debug target polarity to be ready for 1-pass on Sunday DAY
    • Next: Target Work Saturday (8am-3pm)
    • Commission Moller (Dave Gaskell)
    • Access for raster and "easy" target work
    • Ion Calibration for the reference cell
    • N2 Reference Cell Runs
    • MCC wants to do Load Test (30 uA to Hall C)
    • Insert target cell into pick-up coil position, measure polarization
  • Tomorrow
    • Spin Dance during DAY (Dave Gaskell)
    • RF recovery ??
    • Reference cell N2 program
    • Be permissive of access as needed to debug target

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
  • There is no lock available for the ramp rate. This will be fixed Friday morning. Until then, request a more conservative 0.5uA/s ramp rate and carefully monitor.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • The operators have been somewhat careful with our beam orbit locks. Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.

Current Issues

  • Raster "breathing" visible on BPM [3] (not really an issue, just a set of interesting plots).
  • EPR
    • Software operational [4]
    • EPR setup still needs debugging [5]; Suspected solution requires access.

Resolved Issues

  • New alarms:
    • on the HMS polarity using the Hall Probe readout [6]
    • on the status of the insertable half-wave plate [7]
  • SHMS PID trigger tresholds as determined by Simona in Summer were found to be working perfectly for this run as well [8]

Planned Access to the Hall

  • Re-configure carbon foil target Saturday DAY?
    • Needs coordination with RadCon and Techs. Keep in mind Techs only can work ~8am-3pm. Target foils may require RadWorker 2 (hopefully they won't). Need to start as early as possible.
  • Work on Laser RTD #4 (Murchhana & Mingyu, ~1h)
  • Debug EPR setup (Melanie)
  • Check coil polarities
  • Replace bunker fan
  • Check half-way plate status in laser box

Planned Accelerator Configuration Changes

  • Pass-change to 1-pass on Sunday 12/14 DAY
  • Pass-change to 5-pass on Wednesday 12/17 DAY
  • End-of-run (Holiday Break) Friday 12/20
  • Re-startup on 01/10/2020

Special Requests

  • Please be careful with our target :-)